OpenStudy (anonymous):

find the exact value of the remaining trigonometric functions of θ. cosθ=(4)/(5) , 270° < θ < 360°

OpenStudy (rulnick):

theta = acos(4/5) = 36.9 degrees (rounded) so the theta you want is 360-36.9 = 323.1 degrees (rounded)

hero (hero):

not exactly an exact value, but okay

OpenStudy (rulnick):

oh, my mistake. thanks, Hero. I'll get the exact value.

hero (hero):

Yeah, if it hadn't said "find exact value", I would have already posted the solution

OpenStudy (rulnick):

Well, the best we can do is 360 - acos(4/5), or am I missing an obvious closed form for acos(4/5) ... ?